# Projections of standardised energy indices in future climate scenarios

### v1

@@ 1 Introduction
Sources of renewable energy are sensitive to fluctuations in the weather and climate, adding a layer of complexity to this balancing act. Solar and wind energy, for example, are the fastest-growing electricity sources (Cronin et al., 2018; EEA, 2023), and are expected to play a yet more significant role in Europe’s future renewable energy systems. However, both are governed by weather conditions with high spatial and temporal variability, giving these sources an intermittent nature that poses a challenge for consistent energy production (von Bremen, 2010; Raynaud et al., 2018).
@@
Due to the weather-dependence of energy demand and renewable energy production, changes in the climate are expected to affect the characteristics of energy droughts. Recent studies have therefore investigated how renewable energy production will change under future climate conditions (e.g. Yang et al., 2022), and how this will influence the frequency and duration of energy production droughts (Kapica et al., 2024; Zuo et al., 2024). In these studies, energy production droughts are defined as instances where the renewable energy production falls below a pre-determined threshold, corresponding to a quantile of previously observed values. However, the scale of renewable energy production (and energy demand) depends on several factors, such as the region of interest; a daily production of 500 GWh might be abnormally low in one region, but abnormally high in another, depending on the region’s climate and installed capacities. This makes it challenging to compare the characteristics of energy droughts across different regions. For example, how does one systematically assess whether climate change will have a more profound effect on the severity of energy droughts at one region compared to another, if the interpretation of the drought’s severity depends on the region?
Allen and Otero (2023) therefore suggest defining energy droughts in terms of standardised energy indices (see also Stoop et al., 2024). The standardised indices project renewable energy production to a standardised scale, facilitating comparisons between regions with different climates and installed capacities. In this paper, we investigate how the behaviour of standardised energy indices, and the resulting energy droughts, evolve in Europe as a result of climate change. Results are presented for one climate change scenario and several General Circulation Models using CMIP5 data (see Section 3). We obtain similar results to previous studies regarding the frequency and duration of future European energy droughts, but by employing standardised indices, we can additionally make conclusions about the behaviour of standardised energy indices and the magnitude of the corresponding droughts. We further extend previous studies by investigating how installed capacities could be adapted so that energy production droughts do not become more common or severe in the future.
The definition of energy droughts in terms of standardised indices is presented in Section 2. Section 3 presents the data and methods used to study future energy drought characteristics in this work, the results for which are presented in Section 4. A conclusion of the results, and a discussion of the implications of these findings for energy policy and planning, is presented in Section 5. Additional results can be found in the supplementary material.
@@ 2 Energy drought definitions
Following Allen and Otero (2023), we define energy production droughts in terms of the Standardised Renewable Energy Production Index (SREPI). The SREPI is defined as
@@
The time series of past observations used to estimate $F_{P}$ is taken to be daily time series of renewable energy production prior to the current day. This means that the standardised energy indices, and hence the corresponding energy droughts, are defined in terms of the current climate. This is useful when studying properties of energy droughts in the future, since it allows us to investigate how the characteristics of energy droughts will evolve if we maintain the current energy system; we can then deliberately study how specific changes to the energy system, such as increasing installed renewable energy capacities, affect these characteristics in relation to today’s energy system.
Kapica et al. (2024) and Zuo et al. (2024) alternatively study the behaviour of energy droughts in climate change scenarios when droughts are defined as the non-exceedance of a fixed production threshold. If this threshold is defined in terms of quantiles of previous observations, as is commonly the case, then this is equivalent to defining energy droughts using standardised energy indices. The advantage of standardised indices is that the indices are defined on a standardised scale, meaning they can easily be compared for different locations and under different climatic conditions; see Allen and Otero (2023) and Kittel and Schill (2024) for more thorough discussions on different ways to define energy droughts.
Using the definition above, each time step is or is not in an energy drought state. The duration of the drought can be defined as the number of consecutive time steps in a drought state. Similarly, the magnitude of the drought can be quantified by the cumulative SREPI value recorded during the drought (Allen and Otero, 2023); this assumes that the most impactful droughts are those that persist for a long time, even if relatively weak at individual time steps, or when the production is severely lower than expected, even if only for a short time (or both). The standardised index on a given day quantifies the intensity of the drought, and we also consider the average index value over the drought event as a measure of the average drought intensity (Bracken et al., 2024). This provides a convenient framework with which to study the frequency, duration, magnitude, and intensity of drought events.
@@ 3 Methods
To study future energy droughts (EDs), we consider daily data on solar photovoltaic power generation (SPV) and onshore wind power generation (WON) in capacity factor units. Offshore wind generation is not considered due to data availability. The data has been extracted at a sub-country level from the Copernicus Climate Data Store (2024), with “sub-country level” referring to level 2 of the European Union’s NUTS (Nomenclature des Unités Territoriales Statistiques) classification, which corresponds to geographical regions for regional policies.
Following Kapica et al. (2024), we refer to data between 1970 and 2020 as the present day, and data between 2048 and 2098 as the future climate. The analysis is performed using 8 CMIP5 climate models (see Kapica et al., 2024, Table 1 for details) forced by the RCP8.5 scenario. As Bloomfield et al. (2022) remark, identifying and understanding the impacts of future climate conditions on power systems is a significant challenge, primarily due to the multiple types of uncertainties inherent in Global Climate Models. Results are obtained for the mean of each climate model, though, for concision, we often only present the average results across the climate models. We separately analyse EDs in different seasons: winter (DJF), spring (MAM), summer (JJA), and autumn (SON); doing so avoids the EDs clustering in specific seasons with the lowest energy production (see e.g. Allen and Otero, 2023). While new CMIP6 climate scenarios are available, we restrict attention to CMIP5 projections to permit a comparison with previous studies on future EDs (Kapica et al., 2024).
In reality, both wind and solar power are available to energy systems. However, computing total energy production is not possible without access to the installed capacities in each region. Hence, we separately consider wind and solar production droughts, which correspond respectively to abnormally low wind and solar production.These droughts are defined using the recorded wind and solar capacity factors. The total absolute wind (solar) production is equal to the wind (solar) capacity factor multiplied by the installed wind (solar) capacity in the region of interest; hence the capacity factor and production differ by a scaling term. The SREPI is agnostic to the scale of the data, so using capacity factor data will not effect the resulting drought characteristics.
@@
where $\mathrm{CF}_{SPV,t}$ and $\mathrm{CF}_{WON,t}$ are the capacity factors of the photovoltaic and wind energy generators at time $t$, respectively, $\mathrm{CF}_{SPV,max}$ and $\mathrm{CF}_{WON,max}$ are the maximum value of these capacity factors over a reference period, and $0\leq a\leq 1$ represents the proportion of solar power in the solar-wind hybrid generator’s installed capacity. We follow Kapica et al. (2024) by considering this hybrid energy production with $a=0.5$, implying a 50% contribution from SPV and 50% from WON to the energy mix. This choice is discussed further in Section 5. We present some results that illustrate how the behaviour of energy droughts changes for different values of $a$.
@@ 4 Results
4.1 Energy droughts in climate change
@@
In this section, we describe the projected changes in energy droughts for SPV, WON, and HYB under the RCP8.5 emissions scenario. Figure 1 shows the evolution of the SREPI in spring in Berlin, Germany. The SREPI values behave similarly between 1970 and 2020, but are generally much lower in the future climate. The SREPI more frequently attains an extremely low value, falling below -1.28 more often than in the present climate, suggesting an increase in solar energy drought frequency and intensity.
@@
Figure 2 displays time series of annual ED days in Berlin, for all seasons and renewable energy sources. Despite some interannual variability, drought days generally increase in the future climate in all seasons, particularly in winter, with solar production droughts increasing more than wind and hybrid production. The hybrid energy system acts similarly to WON in the colder seasons, and SPV in the warmer seasons. The total number of solar drought days increases by 28% from the present to the future climate.
@@
Figure 1: Time series of SREPI for SPV in one region (Berlin, Germany) during spring. Results are shown for the Model 1 mean in the present (1970-2020) and future (2048-2098) climates. A dashed black line is shown at -1.28.
@@
Figure 2: Time series of moderate energy drought days for one region (Berlin, Germany). Multimodel mean for historical (1970-2020) and RCP8.5 scenario (2048-2098).
@@
Figure 3 displays the spatial distribution of these changes in drought frequency. Solar production droughts are expected to become more common in northern and central Europe, particularly in winter and autumn. The number of winter solar drought days is expected to increase by more than 2.5 per year in Czechia, Poland, Lithuania, Latvia, and Estonia; the colourbar of the plot has been truncated at 2.5. In the colder seasons, changes of more than 2.5 days are expected in northern Sweden. Solar production droughts decrease throughout the year in some regions of Turkey and the Iberian Peninsula. A similar pattern is found when using a more extreme threshold to define EDs (see supplementary material, Section 2).
@@
The projected increase in wind production droughts is markedly lower than for solar. The highest increases in drought frequency occur in Turkey, northern Italy, and western Norway during winter, and in the Scandinavian countries, the UK, and central European countries during summer. In contrast, wind droughts are projected to decrease in summer in southern Europe: wind droughts in Spain, Italy, and Turkey are expected to decrease by more than 2 days per year. Changes are more (less) pronounced in summer (spring). An increase in wind drought frequency over central and northern Europe can be associated with the projected increase in low flow conditions (Otero et al., 2018), though the trend in southern Europe is unclear (e.g. Carvalho et al., 2017; Zhang et al., 2018). We discuss the source of some of these uncertainties in Section 5.
@@
Changes in HYB droughts resemble the changes to SPV droughts, with larger increases in central and northern Europe. HYB droughts are projected to increase less than SPV droughts in winter and autumn (e.g. in Baltic and Scandinavian countries), but increase more than WON droughts in lower latitudes during summer. In spring, the projected increase of HYB droughts is larger than for WON and SPV droughts in northern Europe. This is consistent with HYB following WON during cold seasons and SPV during warm seasons.
@@
Figure 3: Projected changes in moderate energy drought days per year for SPV (a, b, c, d), WON (e, f, g, h) and Hybrid (i, j, k, l). CMIP5 multimodel mean for RCP8.5 scenario (2048-2098). Hatches denote regions of ensemble agreement on the sign of bias; i.e., more than 6 of the ensemble members indicate a bias of the same sign.
@@
In the following, we analyse changes in drought duration and intensity. Figure 4 presents changes in the maximum ED duration in each region. The spatial variability of the maximum drought duration is larger than for the mean duration (see supplementary material, Figure S6), and the climate models only agree on the sign of the change in duration in a few regions. For SPV droughts, an increase in the maximum duration is clear during winter, spring, and autumn in northern Europe: increases of more than 2 days are projected in Scandinavian countries.
@@
Figure 4: Projected changes in max duration of energy drought days for SPV (a, b, c, d), WON (e, f, g, h) and Hybrid (i, j, k, l). Ensemble mean for RCP8.5 scenario (2048-2098). Hatches denote regions of ensemble agreement on the sign of bias; i.e., more than 6 of the ensemble members indicate a bias of the same sign.
@@
These results align with those in Kapica et al. (2024), highlighting that droughts defined using SREPI values exhibit similar properties to those defined using fixed production thresholds. One benefit of employing standardised indices is that we can additionally study the magnitude and intensity of energy droughts across different regions. The magnitude of an ED takes into account both its duration and intensity, and changes in ED magnitude are therefore similar to changes in the mean duration (see Figure S7 of the supplementary material). Figure 5 displays the projected changes in average drought intensity at each region; recall that the average drought intensity is the average SREPI value over the drought event. The intensity is defined on the same scale as the standardised indices. Solar production drought intensity increases in almost all regions and seasons, except in summer, when a decrease in solar drought magnitude is projected in the Iberian Peninsula and some regions in southeast Europe. The HYB energy droughts follow a similar pattern, whereas changes in the average intensity of wind production droughts tend to behave similarly to changes in mean occurrence. For instance, a decrease in intensity is projected at lower latitudes and an increase at higher latitudes; this dipole pattern is more evident in summer.
@@
Changes in ED frequency, duration, and intensity are largest for solar production, particularly over Scandinavian. Since, solar power production is largely determined by global horizontal irradiation and 2m temperature, we explore how these results correspond to projected changes in total cloud cover (supplementary material, Figure S8). There is a clear increase in cloud cover over Scandinavia in all seasons, while a decrease in cloud cover is projected over the Iberian Peninsula for spring, summer, and autumn; this is a region with less increase in SPV energy droughts. This supports the conjecture from Kapica et al. (2024) that a decrease in solar power production can be attributed to more storms and their associated clouds, which decreases solar radiation and therefore increases the likelihood of a solar production drought.
@@
Figure 5: Projected changes in average drought intensity for SPV (a, b, c, d), WON (e, f, g, h) and HYB (i, j, k, l). The average intensity is the average SREPI value across the drought event. Results are shown for the mean across all models in the RCP8.5 scenario (2048-2098). Hatches denote regions where more than 6 of the ensemble members project a change of the same sign.
4.2 Energy systems adaptation to mitigate EDs

### v2

@@ 1 Introduction
Sources of renewable energy are sensitive to fluctuations in the weather and climate, adding a layer of complexity to this balancing act. Solar and wind energy, for example, are the fastest-growing electricity sources (Cronin et al., 2018; EEA, 2023), and are expected to play a yet more significant role in Europe’s future renewable energy systems. However, both are governed by weather conditions with high spatial and temporal variability, giving these sources a variable nature that poses a challenge for consistent energy production (von Bremen, 2010; Raynaud et al., 2018).
@@
Due to the weather-dependence of energy demand and renewable energy production, changes in the climate are expected to affect the characteristics of energy droughts. Recent studies have therefore investigated how renewable energy production will change under future climate conditions (e.g. Yang et al., 2022), and how this will influence the frequency and duration of energy production droughts (Kapica et al., 2024; Zuo et al., 2024). In these studies, energy production droughts are defined as instances where the renewable energy production falls below a pre-determined threshold, corresponding to a quantile of previously observed values. However, the scale of renewable energy production (and energy demand) depends on several factors, such as the region of interest; a daily production of 500 GWh might be abnormally low in one region, but abnormally high in another, depending on the region’s climate and installed capacities. This makes it challenging to compare the characteristics of energy droughts across different regions. For example, how does one systematically assess whether climate change will have a more profound effect on the severity of energy droughts in one region compared to another, if the interpretation of the drought’s severity depends on the region?
Allen and Otero (2023b) therefore suggest defining energy droughts in terms of standardised energy indices (see also Stoop et al., 2024). The standardised indices project renewable energy production to a standardised scale, facilitating comparisons between regions with different climates and installed capacities. In this paper, we investigate how the behaviour of standardised energy indices, and the resulting energy droughts, evolve in Europe as a result of climate change. Results are presented for one climate change scenario and several General Circulation Models using CMIP5 data (see Section 3). We obtain similar results to previous studies regarding the frequency and duration of future European energy droughts, but by employing standardised indices, we can additionally make conclusions about the behaviour of standardised energy indices and the magnitude of the corresponding droughts. We further extend previous studies by investigating how installed capacities could be adapted so that energy production droughts do not become more common or severe in the future.
Due to the available data, we restrict attention to energy production droughts, and do not consider energy supply droughts. As such, the term “energy droughts” henceforth corresponds to energy production droughts. The definition of energy droughts in terms of standardised indices is presented in Section 2. Section 3 presents the data and methods used to study future energy drought characteristics in this work, the results of which are presented in Section 4. These results are summarised in Section 5, before Section 6 discusses the implications of these findings for energy policy and planning, and calls for further work to make more energy data publicly available. Additional results can be found in the supplementary material.
@@ 2 Energy drought definitions
Following Allen and Otero (2023b), we define energy production droughts in terms of the Standardised Renewable Energy Production Index (SREPI). The SREPI is defined as
@@
In this study, the past observations used to estimate $F_{P}$ are daily time series of renewable energy production prior to the current day. This means that the standardised energy indices, and hence the corresponding energy droughts, are defined in terms of the current climate. This is useful when studying properties of energy droughts in the future, since it allows us to investigate how the characteristics of energy droughts will evolve if we maintain the current energy system; we can then deliberately study how specific changes to the energy system, such as increasing installed renewable energy capacities, affect these characteristics in terms of today’s energy system.
Kapica et al. (2024) and Zuo et al. (2024) alternatively study the behaviour of energy droughts in climate change scenarios when droughts are defined as the non-exceedance of a fixed production threshold. If this threshold is defined in terms of quantiles of previous observations, as is commonly the case, then this is equivalent to defining energy droughts using standardised energy indices. The advantage of standardised indices is that the indices are defined on a standardised scale, meaning they can easily be compared for different locations and under different climatic conditions; see Allen and Otero (2023b) and Kittel and Schill (2024) for more thorough discussions on different ways to define energy droughts.
Using the definition above, each time step is or is not in an energy drought state. The duration of the drought can be defined as the number of consecutive time steps in a drought state. Similarly, the magnitude of the drought can be quantified by the cumulative SREPI value recorded during the drought (Allen and Otero, 2023b); this assumes that the most impactful droughts are those that persist for a long time, even if relatively weak at individual time steps, or when the production is severely lower than expected, even if only for a short time (or both). The standardised index on a given day quantifies the intensity of the drought, and we also consider the average index value over the drought event as a measure of the average drought intensity (Bracken et al., 2024). We additionally employ the Compound Drought Magnitude proposed by Bracken et al. (2024), for which we calculate the SREPI separately for wind and solar production, and then average these two indices during drought events; lower values therefore correspond to droughts for which the wind and solar production are both abnormally low. The SREPI therefore provides a convenient framework with which to study the frequency, duration, magnitude, and intensity of drought events. In the following, the SREPI is calculated using the SEI package in R (Allen and Otero, 2023a).
@@ 3 Methods
To study future energy droughts (EDs), we consider daily data on solar photovoltaic power generation (SPV) and onshore wind power generation (WON) in capacity factor units. Offshore wind generation is not considered due to data availability. The data has been extracted at a sub-country level from the Copernicus Climate Data Store (2021) (see also Dubus et al., 2023), with “sub-country level” referring to level 2 of the European Union’s NUTS (Nomenclature des Unités Territoriales Statistiques) classification, which corresponds to geographical regions for regional policies.
Following Kapica et al. (2024), we refer to data between 1970 and 2020 as the present day, and data between 2048 and 2098 as the future climate. The analysis is performed using 8 EURO-CORDEX climate simulations (see Kapica et al., 2024, Table 1 for details) forced by the RCP4.5 and RCP8.5 climate scenarios. Results for the more moderate RCP4.5 scenario are presented in Section 4, while results for the more severe RCP8.5 scenario are provided in the supplementary material (Section 2).
As Bloomfield et al. (2022) remark, identifying and understanding the impacts of future climate conditions on power systems is a significant challenge, primarily due to the multiple types of uncertainties inherent in Global Climate Models. Results are obtained for the mean of each climate model, though, for concision, we often only present the average results across the climate models. We separately analyse energy droughts in different seasons: winter (DJF), spring (MAM), summer (JJA), and autumn (SON); doing so avoids the droughts clustering in specific seasons with the lowest energy production (see e.g. Allen and Otero, 2023b). While new CMIP6 climate scenarios are available, we restrict attention to CMIP5 projections to permit a comparison with previous studies on future energy droughts (Kapica et al., 2024).
We separately consider wind and solar production droughts, which correspond respectively to abnormally low wind and solar production. These droughts are defined using the SREPI as described in Section 2, with the production $P_{t}$ equal to the recorded wind or solar capacity factor at time $t$. The total absolute wind (solar) production is equal to the wind (solar) capacity factor multiplied by the installed wind (solar) capacity in the region of interest; hence the capacity factor and production differ by a scaling term. The SREPI is agnostic to the scale of the data, so using capacity factors rather than total production data will not affect the resulting drought characteristics. Ideally, we would also consider droughts defined using the total renewable energy production from both wind and solar power. However, computing the total energy production is not possible without access to the installed wind and solar capacities in each region. The need for more publicly available energy data is discussed further in Section 6.
@@
where $\mathrm{CF}_{SPV,t}$ and $\mathrm{CF}_{WON,t}$ are the capacity factors of the photovoltaic and wind energy generators at time $t$, respectively, $\mathrm{CF}_{SPV,max}$ and $\mathrm{CF}_{WON,max}$ are the maximum value of these capacity factors over a reference period, and $0\leq a\leq 1$ represents the proportion of solar power in the solar-wind hybrid generator’s installed capacity. Due again to the lack of installed capacity data, this proportion is generally not known. We follow Kapica et al. (2024) by considering this hybrid energy production with $a=0.5$, implying a 50% contribution from SPV and 50% from WON to the energy mix. This choice is discussed further in Section 6, and in Section 4, we present some results that illustrate how the behaviour of energy droughts changes for different values of $a$.
@@ 4 Results
4.1 The frequency of energy droughts
@@
In this section, we describe the projected changes in the frequency of energy droughts for SPV, WON, and HYB under the RCP4.5 emissions scenario. Figure 1 shows the evolution of the SPV SREPI values in spring for the area of Berlin, Germany (the same plot for WON is shown in Figure S5 of the supplementary material). The SREPI behaves similarly between 1970 and 2020, but is generally much lower in the future climate. The SREPI more frequently attains an extremely low value, falling below -1.28 more often than in the present climate, suggesting an increase in solar energy drought frequency and intensity.
@@
Figure 2 displays time series of annual energy drought days in Berlin, for all seasons and renewable energy sources. Despite some interannual variability, drought days generally increase in the future climate in all seasons, particularly in winter, with solar production droughts increasing more than wind and hybrid production. The hybrid energy system acts similarly to WON in the colder seasons, and SPV in the warmer seasons. During winter, the total number of solar drought days increases by 37% from the present to the future climate.
@@
Figure 1: Time series of SREPI for SPV in one region (Berlin, Germany) during spring. Results are shown for the Model 1 mean for historical (1970-2020) and RCP4.5 scenario (2048-2098). A dashed black line is shown at -1.28.
@@
Figure 2: Time series of moderate energy drought days for one region (Berlin, Germany) in different seasons: a) winter, b) spring, c) summer and d) autumn. Multimodel mean for historical (1970-2020) and RCP4.5 scenario (2048-2098).
@@
Figure 3 displays the spatial distribution of these changes in drought frequency. Solar production droughts are expected to become more common in northern and central Europe, particularly in winter and autumn. The number of winter solar drought days is expected to increase by more than 2 per year in Czechia, Poland, Lithuania, Latvia, and Estonia; the colourbar of the plot has been truncated at 2. In the colder seasons, changes of more than 2 days are expected in northern Sweden. Solar production droughts decrease throughout the year in some regions of Turkey and the Iberian Peninsula. A similar pattern is found when using a more extreme threshold to define energy droughts (see Section 3 of the supplementary material).
@@
The projected increase in wind production droughts is markedly lower than for solar. The highest increases in drought frequency occur in Turkey, northern Italy, and western Norway during winter, and in the Scandinavian countries, the UK, and central European countries during summer. In contrast, wind droughts are projected to decrease in summer in southern Europe: wind droughts in Spain, Italy, and Turkey are expected to decrease by more than 2 days per year. Changes are more (less) pronounced in summer (spring). An increase in wind drought frequency over central and northern Europe can be associated with the projected increase in low flow conditions (Otero et al., 2018), though the trend in southern Europe is unclear (e.g. Carvalho et al., 2017; Zhang et al., 2018).
@@
Changes in HYB droughts resemble the changes to SPV droughts, with larger increases in central and northern Europe. HYB droughts are projected to increase less than SPV droughts in winter and autumn (e.g. in Baltic and Scandinavian countries), but increase more than WON droughts in higher latitudes during summer. In spring, the projected increase of HYB droughts is larger than for WON and SPV droughts in northern Europe. This is consistent with HYB following WON during cold seasons and SPV during warm seasons.
@@
Figure 3: Projected changes in moderate energy drought days per year for SPV (a, b, c, d), WON (e, f, g, h) and Hybrid (i, j, k, l). Each column corresponds to a season: a,e,i) winter, b,f,j) spring, c,g,k) summer and d,h,l) autumn. Multimodel mean for RCP4.5 scenario (2048-2098). Hatches denote regions of ensemble agreement on the sign of bias; i.e., more than 6 of the ensemble members indicate a bias of the same sign.
4.2 The duration and intensity of energy droughts
@@
In the following, we analyse changes in energy drought duration and intensity. Figure 4 presents changes in the maximum drought duration in each region. The spatial variability of the maximum drought duration is larger than for the mean duration (see supplementary material, Figure S11), and the climate models only agree on the sign of the change in duration in a few regions. These results should therefore be interpreted with caution. For SPV droughts, an increase in the maximum drought duration is clear during winter, spring, and autumn in northern Europe: increases of more than 2 days are projected in Scandinavian countries.
@@
Figure 4: Projected changes in max duration of energy drought days for SPV (a, b, c, d), WON (e, f, g, h) and Hybrid (i, j, k, l). Each column corresponds to a season: a,e,i) winter, b,f,j) spring, c,g,k) summer and d,h,i) autumn. Multimodel mean for RCP4.5 scenario (2048-2098). Hatches denote regions of ensemble agreement on the sign of bias; i.e., more than 6 of the ensemble members indicate a bias of the same sign.
@@
These results align with those in Kapica et al. (2024), highlighting that droughts defined using SREPI values exhibit similar properties to those defined using fixed production thresholds. One benefit of employing standardised indices is that we can additionally study the magnitude and intensity of energy droughts across different regions. The magnitude of an energy drought takes into account both its duration and intensity, and changes in drought magnitude are therefore similar to changes in the mean duration (see Figure S11 of the supplementary material). Figure 5 displays the projected changes in average drought intensity at each region; recall that the average drought intensity is the average SREPI value over the drought event. The intensity is defined on the same scale as the standardised indices. Solar production drought intensity increases in almost all regions and seasons, except in summer, when a decrease in solar drought magnitude is projected in the Iberian Peninsula and some regions in southeast Europe. The HYB energy droughts follow a similar pattern, whereas changes in the average intensity of wind production droughts tend to behave similarly to changes in mean occurrence. For instance, a decrease in intensity is projected at lower latitudes and an increase at higher latitudes; this dipole pattern is more evident in summer. Following Bracken et al. (2024), we computed the Compound Drought Magnitude, which accounts for simultaneous solar and wind droughts. These results are comparable to the mean intensity of the hybrid generator (Fig. 5i,j,k,l), although they exhibit greater spatial variability and lower model agreement (see Fig. S4 for the historical period and Fig. S13 for the RCP4.5 scenario).
@@
Changes in energy drought frequency, duration, and intensity are largest for solar production, particularly over Scandinavia. Since, solar power production is largely determined by global horizontal irradiation and 2m temperature, we explore how these results correspond to projected changes in total cloud cover (supplementary material, Figure S14). There is a clear increase in cloud cover over Scandinavia in all seasons, while a decrease in cloud cover is projected over the Iberian Peninsula for spring, summer, and autumn; this is a region with less increase in SPV energy droughts. This supports the conjecture from Kapica et al. (2024) that a decrease in solar power production can be attributed to more storms and their associated clouds, which decreases solar radiation and therefore increases the likelihood of a solar production drought.
@@
Figure 5: Projected changes in average drought intensity for SPV (a, b, c, d), WON (e, f, g, h) and HYB (i, j, k, l). Each column corresponds to a season: a,e,i) winter, b,f,j) spring, c,g,k) summer and d,h,i) autumn. The average intensity is the average SREPI value across the drought event. Multimodel mean for RCP4.5 scenario (2048-2098). Hatches denote regions of ensemble agreement on the sign of bias; i.e., more than 6 of the ensemble members indicate a bias of the same sign.
4.3 Energy system adaptation